Robben Manufacturing has the following two possible projects. The required return is 11 percent. Year Project Y Project Z 0 27,200 52.000 21,000 23,000 19,000 21,000 13,200 11,600 14.000 9.600 a. What is the profitability index for each project? (Do not round intermediate calculations and round your answers to 3 decimal places, e.g., 32.161.) Profitability index Project Y Project Z b. What is the NPV for each project? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) NPV Project Y Project Z c. Which, if either, of the projects should the company accept? (Click to select) Neither Project Z Project Y eBook & Resources
Answer #1

Hello Sir/ Mam

PRESENT VALUE SCHEDULE

Y Z
Year Cashflows Present Value Factor Present Value Year Cashflows Present Value
0 -$27,200.00 1.000000 -$27,200.00 0 -$52,000.00 -$52,000.00
1 $13,200.00 0.900901 $11,891.89 1 $21,000.00 $18,918.92
2 $11,600.00 0.811622 $9,414.82 2 $23,000.00 $18,667.32
3 $14,000.00 0.731191 $10,236.68 3 $19,000.00 $13,892.64
4 $9,600.00 0.658731 $6,323.82 4 $21,000.00 $13,833.35
Total $10,667.21 $13,312.22


(a) Profitability Index :

PI
Project Y 1.392
Project Z 1.256

Profitability \ Index=\frac{Present \ Value \ of \ Cash \ Inflows}{Present \ Value \ of \ Cash \ Outflows}

Profitability \ Index(A)=\frac{\$37,867.21 }{\$27,200.00 }=1.392

Profitability \ Index(B)=\frac{\$65,312.22 }{\$52,000.00 }=1.256

(b) NPV : As evident from the Present Value Schedule above,

NPV
Project Y $10,667.21
Project Z $13,312.22

NPV = \sum {Present \ Value \ of \ Cashflows}

(c) THE CORERCT ANSWER IS OPTION B : PROJECT Z

Sir/ Mam, in case of multiple questions asked in a single post, we are required just to answer the first. Hence, I'll request you to please ask the second one in a seperate post. Thanks!

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.